C = 60 + 0.75Td I = 79 G = 45 T = 0.2Y - 25 Yd = Y - T Y = C + I + G What is the quilibrium income:
Answer #1

We have the following information

Consumption: C = 60 + 0.75Yd

Investment: I = 79

Government Spending: G = 45

Taxes: T = 0.2Y – 25

Y = Income

Yd = Y – T (given)

Y = C + I + G

Y = 60 + 0.75Yd + 79 + 45

Y = 60 + 0.75(Y – T) + 124

Y = 60 + 0.75(Y – 0.2Y + 25) + 124

Y = 60 + 0.75(0.8Y + 25) + 124

Y = 60 + 0.6Y + 18.75 + 124

Y = 0.6Y + 202.75

Y – 0.6Y = 202.75

0.4Y = 202.75

Equilibrium Income = 506.89